# Physics-Informed Time-Integrated DeepONet: Temporal Tangent Space Operator Learning for High-Accuracy Inference

### v2

@@ Abstract
Accurately modeling and inferring solutions to time-dependent partial differential equations (PDEs) over extended temporal horizons remains a core challenge in scientific machine learning. Traditional full rollout (FR) methods, predicting entire trajectories in a single pass, often fail to capture the causal dependencies inherent to dynamical systems and exhibit poor generalization outside the training time horizon. In contrast, autoregressive (AR) approaches, which evolve the system step by step, are prone to error accumulation, as predictions at each time step depend on potentially erroneous prior outputs. These shortcomings limit the long-term accuracy and reliability of both strategies. To overcome these issues, we introduce Physics-Informed Time-Integrated Deep Operator Network (PITI-DeepONet), an operator learning framework designed for stable and accurate long-term time evolution, well beyond the training time horizon. PITI-DeepONet employs a dual-output DeepONet architecture trained via either fully physics-informed or hybrid physics- and data-driven objectives. The training enforces consistency between the learned temporal derivative and its counterpart obtained via automatic differentiation. Rather than directly forecasting future states, the network learns the time-derivative operator from the current state, which is then integrated using classical time-stepping schemes - such as explicit Euler, fourth-order Runge-Kutta, or second-order Adams-Bashforth-Moulton - to advance the solution sequentially in time. Additionally, the framework can leverage residual monitoring during inference to estimate prediction quality and detect when the system transitions outside the training domain. Applied to benchmark problems, PITI-DeepONet demonstrates enhanced accuracy and stability over extended inference time horizons when compared to traditional methods. Mean relative $\mathcal{L}_{2}$ errors reduced by 84% (versus FR) and 79% (versus AR) for the one-dimensional heat equation; by 87% (versus FR) and 98% (versus AR) for the one-dimensional Burgers equation; and by 42% (versus FR) and 89% (versus AR) for the two-dimensional Allen-Cahn equation. By moving beyond classic FR and AR schemes, PITI-DeepONet paves the way for more reliable, long-term integration of complex, time-dependent PDEs.
@@ 1 Introduction
Neural operators (NOs) have emerged as powerful surrogate models in recent years, achieving remarkable success in both forward and inverse problems  [lu2021learning, li2021fourier, Kobayashi2024, kag2024learning]. These include applications in additive manufacturing [Liu2024, Kushwaha2024], nuclear energy systems [Kobayashi2024b], and biomedical topics [goswami2022neural] to name a few. Physics-informed approaches [Karniadakis2021, Wang2021, Goswami2023] have also gained prominence, enabling the training process to dispense entirely, or in part, with labeled data pairs by relying instead on physical laws and residual minimization, akin to solving classical initial-boundary value problems of partial differential equations (PDEs). However, the way operator learning is employed differs fundamentally from traditional numerical methods for solving differential equations. Common approaches include full rollout (FR) models, which train on an entire spatiotemporal domain and then perform inference on that same domain [lin2023learning, wang2023long].
@@
While the FR approach avoids treating time evolution explicitly, it often struggles with extrapolation beyond the training domain and fails to leverage the Markovian nature of many dynamical systems, where subsequent states inherently depend on previous ones. Alternatively, discrete autoregressive rollout (AR) methods have been introduced but face the limitations of training domain generalization and an additional susceptibility to error accumulation in time-dependent systems [mccabe2023stabilityautoregressiveneuraloperators]. As a result, both approaches are ill-suited for making reliable long-term predictions akin to those obtained via classical simulations, limiting their utility as surrogate models.
@@
Several recent efforts have sought to overcome these limitations by introducing architectural modifications in the NO to provide long-time prediction capabilities. For instance, in  [diab2025temporal], an additional network was introduced to better capture temporal dependencies, while NO outputs were processed further by augmenting recurrent architectures to improve sequence modeling [michalowska2024neural]. Other approaches introduce memory mechanisms [buitrago2024benefits, he2024sequential] to mitigate the non-Markovian behavior inherent in rollout-based methods. Nonetheless, these methods predominantly adopt generic time-series solutions without exploiting the rich dynamical information present in time-dependent systems.
@@
The time-integration embedded operator learning framework introduced [2024Nayak-TI-DeepONet] addresses these limitations by leveraging the system’s intrinsic dynamical structure, enabling near real-time inference, and shifting part of the computational load to an offline phase. This framework learns the discrete tangent from data and reuses it during inference through a numerical time integrator. Building on this foundation, we propose a significant advancement: a continuous temporal tangent space operator, which generalizes the notion of learned time evolution beyond discrete updates, transitioning the training process from a purely data-driven approach to a combination of physics-informed and data-driven methodologies. Additionally, we incorporate a residual prediction mechanism based on the current system state, which serves as an effective error predictor. This online residual tracking provides a means to assess prediction quality with minimal computational overhead, offering new opportunities for robust and efficient operator learning.
@@ 2.1 Time-Integrator embedded Deep Operator Network
The Time-Integrator embedded Deep Operator Network (TI-DeepONet) framework [2024Nayak-TI-DeepONet] transforms the task of temporal prediction in dynamical systems into a derivative learning problem. Rather than predicting future states in an autoregressive manner, the model learns a discrete approximation of the system’s temporal tangent space. Concretely, the current system state $\mathbf{u}^{n}$ is passed through a DeepONet [lu2021learning], which outputs the estimated temporal derivative $\hat{\mathbf{u}}_{t}^{n}$. This derivative is then integrated numerically via schemes like Euler or Runge-Kutta to yield the predicted next state $\hat{\mathbf{u}}^{n+1}$. The model is trained using a supervised loss computed between $\hat{\mathbf{u}}^{n+1}$ and the true state $\mathbf{u}^{n+1}$ and learns an approximate tangent between the data pairs via the underlying scheme as a discrete operator. This setup tightly couples operator learning with time integration and provides a modular structure, where physical priors and numerical schemes can be systematically incorporated.
@@ 2.2 Physics-Informed Time-Integrated Deep Operator Network
In the following, we provide a comprehensive overview of how the physics-informed continuous tangent space operator is derived, along with a discussion on the training process, data sampling, and inference setup.
@@ 2.2.1 Network and Training Setup
To learn the temporal tangent ($\mathbf{u}_{t}^{n}$), we use a physics-informed DeepONet that performs the mapping $\mathcal{G}:\mathbf{u}^{n}\rightarrow{\hat{\mathbf{u}}}^{n},\widetilde{{\mathbf{u}}}^{n}_{t}$, where $\mathbf{u}^{n}$ is the input field, $\hat{\mathbf{u}}^{n}$ is the reconstructed input field, and ${\widetilde{\mathbf{u}}}^{n}_{t}$ is the predicted temporal tangent. For physics-informed training, we use automatic differentiation (AD) on $\hat{\mathbf{u}}^{n}$ to obtain the derivatives necessary to compute the loss terms: (1) PDE residual, $\mathcal{L}_{\rm{PDE}}$, (2) initial condition (IC) loss, $\mathcal{L}_{\rm{IC}}$, and (3) boundary condition (BC) loss, $\mathcal{L}_{\rm{BC}}$. This part is identical to physics-informed training of any operator [Wang2021]. In addition, we introduce two new losses as follows. First, the discrepancy between $\mathbf{u}^{n}$ and $\hat{\mathbf{u}}^{n}$ is quantified as reconstruction loss $\mathcal{L}_{\rm{R}}$. Second, the consistency of the temporal derivative with the field is ensured using a consistency loss, $\mathcal{L}_{\mathrm{C}}$, between the network-predicted derivative, $\widetilde{\mathbf{u}}_{t}^{n}$, and the temporal derivative obtained via automatic differentiation (AD) on $\hat{\mathbf{u}}^{n}$, i.e., $\hat{\mathbf{u}}_{t}^{n}$. To accommodate the temporal dimension in this formulation, phantom values (e.g., zeros) are included as placeholders in the time input to the trunk network (to indicate the current time), alongside other coordinate inputs. Nonetheless, this still effectively reduces the dimensionality of the underlying problem to some extent as only a single value for the time domain has to be sampled. A schematic of the architecture described above is presented in Fig. 1. Moreover, as shown in the schematic, data-driven losses can be additionally imposed on the network outputs to achieve a hybrid setup. In total, this leads to 6 loss terms that are aggregated as the total loss. Note that all losses are computed as mean squared errors (MSE) and then combined with weighting factors $\mathbf{\lambda}=\{\lambda_{\rm{PDE}},\lambda_{\rm{R}},\lambda_{\rm{BC}},\lambda_{\rm{C}},\lambda_{\rm{u}},\lambda_{\rm{u_{t}}}\}$. The total loss reads $\mathcal{L}=\lambda_{\rm{PDE}}\mathcal{L}_{\rm{PDE}}+\lambda_{\rm{R}}\mathcal{L}_{\rm{R}}+\lambda_{\rm{BC}}\mathcal{L}_{\rm{BC}}+\lambda_{\rm{C}}\mathcal{L}_{\rm{C}}+\lambda_{\rm{u}}\mathcal{L}_{\rm{u}}+\lambda_{\rm{u_{t}}}\mathcal{L}_{\rm{u_{t}}}$, with the latter two being the data-driven losses, which are optional. Furthermore, the reconstruction loss and the data-driven loss for $\mathbf{u}^{n}$ have the same formal structure, but are assigned different weights within the loss function to balance the influence of the data-driven and physics-informed terms. This allows for a purely physics-informed or a hybrid setup. All models are optimized using ADAM [Adam2014] with an exponential decay schedule for the learning rate. Note that there are other ways of formulating the physics-informed TI-DeepONet, but each have their disadvantages compared to the proposed method: i) one could discard the output, ${\widetilde{\mathbf{u}}}^{n}_{t}$, and employ only AD to obtain the time derivative in inference. However, this would significantly increase the computational demand and reduce inference speed; ii) one could skip the consistency loss and directly incorporate the output, ${\widetilde{\mathbf{u}}}^{n}_{t}$, in the PDE residual. However, as will be shown later in this work, the resulting network output is comparable or even inferior to our setup, so the more general formulation is preserved.
@@ 2.2.2 Sampling
Key to learning a temporal tangent space operator is to span the domain of solution states during inference with the limited samples used during training. Unlike other methods, which typically involve time evolution as seen in the traditional use of operators, this approach focuses solely on the time derivative of the current state. Consequently, a more meticulous sampling strategy is required. One potential remedy is to increase both the number and diversity of randomly sampled ICs to effectively span the relevant solution manifold. However, achieving sufficient coverage of the solution space through random sampling alone can be daunting—especially in the absence of domain knowledge to guide and reduce the training data requirements. In this work, we adopt a different strategy that leverages available domain knowledge in the form of solution samples $\mathbf{u}^{n}$ and corresponding time derivatives ${\mathbf{u}}^{n}_{t}$ at fixed time instances. These samples may originate from experiments, analytical solutions, numerical simulations, or even pre-trained surrogate models operating over longer time horizons. Once obtained, they are aggregated and used as inputs to the branch network, treated equivalently to IC samples. The trunk network, in turn, receives these inputs with phantom zeroes in the time dimension, omitting explicit time encoding, alongside spatial coordinates and any additional parameters.
@@ 2.2.3 Inference Setup
Having learned a continuous temporal tangent space operator, we can now exploit this for time-stepping-based inference. For each available state as branch input, the network predicts a replica (or reconstruction) of the input state as well as the corresponding temporal tangent (or time-derivative) of that state. To advance the solution in time, we utilize well-established, stable, and accurate time-stepping schemes commonly used for solving ordinary differential equations (ODEs). There are no constraints imposed on parameters such as model complexity, choice of numerical method, number of iterations, or time step size. To illustrate the flexibility of our approach, three representative numerical time integration schemes are considered.
@@
Another notable property of this method is the ability to monitor the residual, $\mathcal{R}(\mathbf{u}^{n},\hat{\mathbf{u}}^{n})=(\mathbf{u}^{n}-\hat{\mathbf{u}}^{n})^{2}$, computed using the input state $\mathbf{u}^{n}$ and the network prediction $\hat{\mathbf{u}}^{n}$. In theory, a well-learned state should correspond to a small residual, indicating consistency between the model’s prediction and the governing dynamics. Conversely, a large residual may suggest that the current input lies outside the distribution of the training data, potentially leading to inaccurate predictions by the learned operator. This behavior effectively provides an intrinsic estimate of prediction quality, enabling a form of self-assessment within the inference process.

### v3

# arXiv:2508.05190v3[cs.LG] 27 Feb 2026

@@ Abstract
Accurately modeling and inferring solutions to time-dependent partial differential equations (PDEs) over extended temporal horizons remains a core challenge in scientific machine learning. Traditional full rollout (FR) methods, predicting entire trajectories in a single pass, often fail to capture the causal dependencies inherent to dynamical systems and exhibit poor generalization outside the training time horizon. In contrast, autoregressive (AR) approaches, which evolve the system step by step, are prone to error accumulation, as predictions at each time step depend on potentially erroneous prior outputs. These shortcomings limit the long-term accuracy and reliability of both strategies. To overcome these issues, we introduce Physics-Informed Time-Integrated Deep Operator Network (PITI-DeepONet), an operator learning framework designed for stable and accurate long-term time evolution, well beyond the training time horizon. PITI-DeepONet employs a dual-output DeepONet architecture trained via either fully physics-informed or hybrid physics- and data-driven objectives. The training enforces consistency between the learned temporal derivative and its counterpart obtained via automatic differentiation. Rather than directly forecasting future states, the network learns the time-derivative operator from the current state, which is then integrated using classical time-stepping schemes - such as explicit Euler, fourth-order Runge-Kutta, second-order Adams-Bashforth-Moulton, or implicit Euler - to advance the solution sequentially in time. Additionally, the framework supports residual monitoring during inference to estimate prediction quality and flags when the learned temporal tangent becomes unreliable, e.g., outside the training domain. Applied to benchmark problems, PITI-DeepONet demonstrates enhanced accuracy and stability over extended inference time horizons when compared to traditional methods. Mean relative $\mathcal{L}_{2}$ errors reduced by 84% (versus FR) and 79% (versus AR) for the one-dimensional heat equation; by 87% (versus FR) and 98% (versus AR) for the one-dimensional Burgers equation; by 42% (versus FR) and 89% (versus AR) for the two-dimensional Allen-Cahn equation; and by 58% (vs. FR) and 61% (vs. AR) for the one-dimensional Kuramoto-Sivashinsky equation. By moving beyond classic FR and AR schemes, PITI-DeepONet paves the way for more reliable, long-term integration of complex, time-dependent PDEs.
@@ 1 Introduction
Neural operators (NOs) have emerged as powerful surrogate models in recent years, achieving remarkable success in both forward and inverse problems  Lu et al. [2021]; Li et al. [2021]; Kobayashi et al. [2024]; Kag et al. [2024]. These include applications in additive manufacturing Liu et al. [2024]; Kushwaha et al. [2024], nuclear energy systems Kobayashi and Alam [2024], and biomedical topics Goswami et al. [2022] to name a few. Physics-informed approaches Karniadakis et al. [2021]; Wang et al. [2021]; Goswami et al. [2023] have also gained prominence, enabling the training process to dispense entirely, or in part, with labeled data pairs by relying instead on physical laws and residual minimization, akin to solving classical initial-boundary value problems of partial differential equations (PDEs). However, the way operator learning is employed differs fundamentally from traditional numerical methods for solving differential equations. Common approaches include full rollout (FR) models, which train on an entire spatiotemporal domain and then perform inference on that same domain Lin et al. [2023]; Wang and Perdikaris [2023].
@@
While the FR approach avoids treating time evolution explicitly, it often struggles with extrapolation beyond the training domain and fails to leverage the Markovian nature of many dynamical systems, where subsequent states inherently depend on previous ones. Alternatively, discrete autoregressive rollout (AR) methods have been introduced but face the limitations of training domain generalization and an additional susceptibility to error accumulation in time-dependent systems McCabe et al. [2023]. As a result, both approaches are ill-suited for making reliable long-term predictions akin to those obtained via classical simulations, limiting their utility as surrogate models.
@@
Several recent efforts have sought to overcome these limitations by introducing architectural modifications in the NO to provide long-time prediction capabilities. For instance, in  Diab and Al-Kobaisi [2025], an additional network was introduced to better capture temporal dependencies, while NO outputs were processed further by augmenting recurrent architectures to improve sequence modeling Michałowska et al. [2024]. Other approaches introduce memory mechanisms Buitrago et al. [2024]; He et al. [2024] to mitigate the non-Markovian behavior inherent in rollout-based methods. Nonetheless, these methods predominantly adopt generic time-series solutions without exploiting the rich dynamical information present in time-dependent systems.
@@
The time-integration embedded operator learning framework introduced Nayak and Goswami [2025] addresses these limitations by leveraging the system’s intrinsic dynamical structure, enabling near real-time inference, and shifting part of the computational load to an offline phase. This framework learns the discrete tangent from data and reuses it during inference through a numerical time integrator. Building on this foundation, we propose a significant advancement: a continuous temporal tangent space operator, which generalizes the notion of learned time evolution beyond discrete updates, transitioning the training process from a purely data-driven approach to a combination of physics-informed and data-driven methodologies. Additionally, we incorporate a residual prediction mechanism based on the current system state, which serves as an effective error predictor. This online residual tracking provides a means to assess prediction quality with minimal computational overhead, offering new opportunities for robust and efficient operator learning.
@@ 2.1 Time-Integrator embedded Deep Operator Network
The Time-Integrator embedded Deep Operator Network (TI-DeepONet) framework Nayak and Goswami [2025] transforms the task of temporal prediction in dynamical systems into a derivative learning problem. Rather than predicting future states in an autoregressive manner, the model learns a discrete approximation of the system’s temporal tangent space. Concretely, the current system state $\mathbf{u}^{n}$ is passed through a DeepONet Lu et al. [2021], which outputs the estimated temporal derivative $\hat{\mathbf{u}}_{t}^{n}$. This derivative is then integrated numerically via schemes like Euler or Runge-Kutta to yield the predicted next state $\hat{\mathbf{u}}^{n+1}$. The model is trained using a supervised loss computed between $\hat{\mathbf{u}}^{n+1}$ and the true state $\mathbf{u}^{n+1}$ and learns an approximate tangent between the data pairs via the underlying scheme as a discrete operator. This setup tightly couples operator learning with time integration and provides a modular structure, where physical priors and numerical schemes can be systematically incorporated.
@@ 2.2 Physics-Informed Time-Integrated Deep Operator Network
In contrast, the proposed physics-informed time-integrated deep operator network (PITI-DeepONet) employs a physics-informed ansatz to learn a continuous temporal tangent operator that maps the current state directly to its time derivative. While TI-DeepONet learns a discrete, fixed-$\Delta t$ tangent map that is tied to the training time step and uses learnable coefficients to compensate when integrating with a chosen numerical scheme, PITI-DeepONet decouples operator learning from time integration: once the continuous tangent is learned, inference can use any standard time integrator with arbitrary step sizes $\Delta t$ without requiring learnable integration weights. In the following, we provide a comprehensive overview of how the physics-informed continuous tangent space operator is derived, along with a discussion on the training process, data sampling, and inference setup. We want to emphasize that PITI is an ansatz-level construction and therefore operator-agnostic, i.e., it can be combined with any neural operator backbone as long as a physics-informed instantiation of that operator is feasible.
@@ 2.2.1 Network and Training Setup
To learn the temporal tangent ($\mathbf{u}_{t}^{n}$), we use a physics-informed DeepONet that performs the mapping $\mathcal{G}:\mathbf{u}^{n}\rightarrow{\hat{\mathbf{u}}}^{n},\widetilde{{\mathbf{u}}}^{n}_{t}$, where $\mathbf{u}^{n}$ is the input field, $\hat{\mathbf{u}}^{n}$ is the reconstructed input field, and ${\widetilde{\mathbf{u}}}^{n}_{t}$ is the predicted temporal tangent. For physics-informed training, we use automatic differentiation (AD) on $\hat{\mathbf{u}}^{n}$ to obtain the derivatives necessary to compute the loss terms: (1) PDE residual, $\mathcal{L}_{\rm{PDE}}$, (2) initial condition (IC) loss, $\mathcal{L}_{\rm{IC}}$, and (3) boundary condition (BC) loss, $\mathcal{L}_{\rm{BC}}$. This part is identical to physics-informed training of any operator Wang et al. [2021]. In addition, we introduce two new losses as follows. First, the discrepancy between $\mathbf{u}^{n}$ and $\hat{\mathbf{u}}^{n}$ is quantified as reconstruction loss $\mathcal{L}_{\rm{R}}$. Second, the consistency of the temporal derivative with the field is ensured using a consistency loss, $\mathcal{L}_{\mathrm{C}}$, between the network-predicted derivative, $\widetilde{\mathbf{u}}_{t}^{n}$, and the temporal derivative obtained via automatic differentiation (AD) on $\hat{\mathbf{u}}^{n}$, i.e., $\hat{\mathbf{u}}_{t}^{n}$. To accommodate the temporal dimension in this formulation, phantom values (e.g., zeros) are included as placeholders in the time input to the trunk network (to indicate the current time), alongside other coordinate inputs. Nonetheless, this still effectively reduces the dimensionality of the underlying problem to some extent as only a single value for the time domain has to be sampled. A schematic of the architecture described above is presented in Fig. 1. Moreover, as shown in the schematic, data-driven losses can be additionally imposed on the network outputs to achieve a hybrid setup. Importantly, PITI does not inherently require time-derivative ground truth: the data-driven term $\mathcal{L}_{\rm{u_{t}}}$ is optional and is used to stabilize and speed up training in this implementation, while ground-truth $\mathbf{u}_{t}$ is additionally required to quantify temporal-tangent errors on held-out test data. In our benchmarks, $\mathbf{u}_{t}$ targets are obtained by evaluating the known PDE right-hand side at stored snapshots (no additional PDE solves; for explicit schemes, this quantity is already computed during time stepping). Therefore, the additional computational cost is minimal compared to generating the reference solution trajectories, which AR and FR training also require. In total, this leads to 6 loss terms that are aggregated as the total loss. Note that all losses are computed as mean squared errors (MSE) and then combined with weighting factors $\mathbf{\lambda}=\{\lambda_{\rm{PDE}},\lambda_{\rm{R}},\lambda_{\rm{BC}},\lambda_{\rm{C}},\lambda_{\rm{u}},\lambda_{\rm{u_{t}}}\}$. The total loss reads $\mathcal{L}=\lambda_{\rm{PDE}}\mathcal{L}_{\rm{PDE}}+\lambda_{\rm{R}}\mathcal{L}_{\rm{R}}+\lambda_{\rm{BC}}\mathcal{L}_{\rm{BC}}+\lambda_{\rm{C}}\mathcal{L}_{\rm{C}}+\lambda_{\rm{u}}\mathcal{L}_{\rm{u}}+\lambda_{\rm{u_{t}}}\mathcal{L}_{\rm{u_{t}}}$, with the latter two being the data-driven losses, which are optional. Furthermore, the reconstruction loss and the data-driven loss for $\mathbf{u}^{n}$ have the same formal structure, but are assigned different weights within the loss function to balance the influence of the data-driven and physics-informed terms. This allows for a purely physics-informed or a hybrid setup. All models are optimized using ADAM Kingma and Ba [2017] with an exponential decay schedule for the learning rate. Note that there are other ways of formulating the PITI-DeepONet, but each have their disadvantages compared to the proposed method: i) one could discard the output, ${\widetilde{\mathbf{u}}}^{n}_{t}$, and employ only AD to obtain the time derivative in inference. However, this would significantly increase the computational demand and reduce inference speed; ii) one could skip the consistency loss and directly incorporate the output, ${\widetilde{\mathbf{u}}}^{n}_{t}$, in the PDE residual. This becomes the special formulation for explicit time-evolution PDEs of the form $u_{t}=\mathcal{N}(\cdot)$, where the time derivative appears in isolation. In this case, we can form the PDE residual by inserting the predicted tangent $\widetilde{\mathbf{u}}_{t}^{n}$ directly into the PDE loss, which avoids any automatic differentiation with respect to $t$ and reduces the need for the consistency loss. Consequently, the trunk does not require an explicit time input, and no dummy 0 placeholders are needed. Nevertheless, as will be shown later in this work, the resulting network output is comparable or even inferior to our setup, so the more general formulation is preserved.
@@ 2.2.2 Sampling
Key to learning a temporal tangent space operator is to span the domain of solution states during inference with the limited samples used during training. Unlike other methods, which typically involve time evolution as seen in the traditional use of operators, this approach focuses solely on the time derivative of the current state. Consequently, a more meticulous sampling strategy is required. One potential remedy is to increase both the number and diversity of randomly sampled ICs to effectively span the relevant solution manifold. However, achieving sufficient coverage of the solution space through random sampling alone can be daunting—especially in the absence of domain knowledge to guide and reduce the training data requirements. In this work, we adopt a different strategy that leverages available domain knowledge in the form of solution samples $\mathbf{u}^{n}$ and corresponding time derivatives ${\mathbf{u}}^{n}_{t}$ at fixed time instances. These samples may originate from experiments, analytical solutions, numerical simulations, or even pre-trained surrogate models operating over longer time horizons. Once obtained, they are aggregated and used as inputs to the branch network, treated equivalently to IC samples. The trunk network, in turn, receives these inputs with phantom zeroes in the time dimension, omitting explicit time encoding, alongside spatial coordinates and any additional parameters. As PITI is queried at the current state during rollout, ensuring reliable performance necessitates that the training branch inputs encompass the state distribution anticipated at inference. Therefore, the sampling strategy cannot be completely agnostic: it needs to be constructed using information specific to the problem (such as expected IC and parameter ranges) to guarantee that the learned temporal tangent is accurate on the states observed during inference, instead of targeting a single universal operator applicable to any arbitrary unseen regimes.
@@ 2.2.3 Inference Setup
Having learned a continuous temporal tangent space operator, we can now exploit this for time-stepping-based inference. For each available state as branch input, the network predicts a replica (or reconstruction) of the input state as well as the corresponding temporal tangent (or time-derivative) of that state. To advance the solution in time, we utilize well-established, stable, and accurate time-stepping schemes commonly used for solving ordinary differential equations (ODEs). There are no constraints imposed on parameters such as model complexity, choice of numerical method, number of iterations, or time step size. Since PITI learns a continuous temporal tangent, the subsequent time integration is decoupled from the training setup and can be performed with any numerical scheme during inference, with accuracy governed by the integrator’s standard stability and order properties. To illustrate the flexibility of our approach, the following representative numerical time integration schemes are considered.
@@
Lastly, the backward Euler method (BE) is included as an A-stable implicit method. The initial guess is formed using explicit Euler, and the next guesses are computed by solving a non-linear fixed-point relation with damped Picard iterations:
$\text{Initialization:}\quad\mathbf{u}^{n+1,(0)}$ $=\mathbf{u}^{n}+\Delta t\,\mathcal{G}_{\boldsymbol{\theta}}\!\left(\mathbf{u}^{n}\right),$ (9)
$\text{Picard map:}\quad\widetilde{\mathbf{u}}^{n+1,(k+1)}$ $=\mathbf{u}^{n}+\Delta t\,\mathcal{G}_{\boldsymbol{\theta}}\!\left(\mathbf{u}^{n+1,(k)}\right),$ (10)
$\text{Relaxation:}\quad\mathbf{u}^{n+1,(k+1)}$ $=(1-\omega)\,\mathbf{u}^{n+1,(k)}+\omega\,\widetilde{\mathbf{u}}^{n+1,(k+1)},\qquad k=0,\dots,K-1.$ (11)
where $\omega\in[0,1]$ is a tunable damping parameter and $k$ the fixed number of iterations. Within this work, we use $K=16$ iterations with $\omega=0.4$.
@@
Another notable property of this method is the ability to monitor the residual, $\mathcal{R}(\mathbf{u}^{n},\hat{\mathbf{u}}^{n})=(\mathbf{u}^{n}-\hat{\mathbf{u}}^{n})^{2}$, computed using the input state $\mathbf{u}^{n}$ and the network prediction $\hat{\mathbf{u}}^{n}$. In theory, a well-learned state should correspond to a small residual, indicating consistency between the model’s prediction and the governing dynamics. Conversely, a large residual may suggest that the current input lies outside the distribution of the training data, potentially leading to inaccurate predictions by the learned operator. This behavior effectively provides an intrinsic estimate of prediction quality, enabling a form of self-assessment within the inference process. As a large residual indicates that the learned time derivative may be unreliable, continuing the rollout would propagate incorrect derivative inputs, and a practical deployment could switch to an alternative derivative source, e.g., a different surrogate or a conventional solver. In this work, we use residual monitoring solely for reporting and do not adapt the inference procedure based on $\mathcal{R}$.